OpenStudy (anonymous):

Write the equation of each line in slope intercept form, point slope form, and standard form: Slope of 3 and passes through (-3,5)

OpenStudy (mathstudent55):

Slope-intercept form is y = mx + b, where m = slope, and b = y-intercept, or point (0, b). Point-slope form is y - y1 = m(x - x1), where m = slope, and the point is (x1, y1). Standard form is ax + by = c. Since you are given the slope and a point, start with slope-point form: y - y1 = m(x - x1) y - 5 = 3[ x - (- 3)] y - 5 = 3(x + 3) Now you can solve for y, and get the slope-intercept form. After that, you can get it in standard form, ax + by = c